HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025, Page 1 of 6  Not logged in ELOG logo
ID Date Author Category Subject Text Attachments
  108   Thu Apr 3 03:48:40 2025 JEL, ZNRunsScreenshots
The peaks in the analyzer are slowly shifting
back to the original position. 
 2025-04-03_03-47-28-530.png2025-04-03_03-47-11-337.png2025-04-03_03-47-42-665.png 
  107   Thu Apr 3 02:59:15 2025 JEL, ZNRunsScreenshot, beam received, baem lost, beam back againBeam is back for 10 mins-

The peaks on the analyzer appear shifted
 2025-04-03_02-58-06-109.png2025-04-03_02-57-54-019.png2025-04-03_02-57-31-430.png 
  106   Thu Apr 3 02:52:27 2025 Zac, JohanGeneralBeam is back at 02:52    
  105   Thu Apr 3 02:01:29 2025 Zac, JohanGeneralQuadrupole issuesHKR has called someone to fix the quadrupole,
they think it will take about an hour.
  
  104   Thu Apr 3 01:56:57 2025 Zac, JohanGeneralProblems with one of the quadrupolesSince around 01:45 we get spills but see nothing
on analysers. HKR says there is a problem
with one of the quadrupoles; they are working
  
  103   Thu Apr 3 01:29:49 2025 Zac, JohanGeneralBeam is back at 01:27    
  102   Thu Apr 3 01:19:44 2025 Zac, JohanGeneralBeam fluctuationsBeam gone again between 01:14 and 01:18, then
again at 01:19. HKR says there are problems
with the source.
  
  101   Thu Apr 3 01:14:16 2025 Zac, JohanGeneralBeam is back at 01:12    
  100   Thu Apr 3 00:49:00 2025 Zac, JohanGeneralBeam lost at 00:46    
  99   Wed Apr 2 23:40:23 2025 JEL, ZNGeneralScreenshots taken just as beam came back
We are seeing spills being received but no
peaks are seen in the analyzers. 
 2025-04-02_23-39-29-655.png2025-04-02_23-37-10-651.png2025-04-02_23-36-58-664.png 
  98   Wed Apr 2 23:33:43 2025 ZacAcceleratorBeam is gone at 23:29    
  97   Wed Apr 2 23:27:35 2025 Zac, JohanAcceleratorBeam is back 23:25    
  96   Wed Apr 2 23:15:52 2025 Zac, JohanAcceleratorNo injections from SISNight shift starts with no injections from
SIS! We still have triggers; we called HKR
and are waiting for response.
  
  95   Wed Apr 2 22:47:25 2025 ID, HA (Radio Shahab as background)GeneralLassie and Schottky TV screentshotsWe managed to take some screenshots while
Shahab harrassed us with AI-created "RammstAIn"
music... happy to end my shift now!
 2025-04-02_22-44-18-420.png2025-04-02_22-45-12-849.png2025-04-02_22-55-14-900.png 
  94   Wed Apr 2 21:22:41 2025 ID, HAGeneralLassie Come-Home Screenshots (wuff!)(Who let the dog out?)
Sharing with FRS 6:1 since a while now.
 2025-04-02_21-25-00-716.png2025-04-02_21-25-42-535.png2025-04-02_21-26-12-022.png 
  93   Wed Apr 2 20:57:34 2025 ShahabDetectors10GB cables for the DAQ rackWe have been working together with IT on the
speed of data transfer: with the new 10 GB
interface we can now transfer data from NAS
  
  92   Wed Apr 2 20:15:29 2025 ID, HAGeneralLassie screenshots + Schottky TVScreeeeeeeenshooooots!  2025-04-02_20-15-48-827_tcl1064.png2025-04-02_20-17-14-017.png2025-04-02_20-18-29-150.png 
  91   Wed Apr 2 19:13:50 2025 ID, HAGeneralLassie screenshots and Schottky screenshotLooks a bit more regular now (FRS people having
dinner???)
 2025-04-02_19-10-29-388_tcl1064.png2025-04-02_19-14-53-428_tcl1146.png2025-04-02_19-19-23-765.png 
  90   Wed Apr 2 18:31:52 2025 IDGeneralFYI for morning shifts ThuThe PLEIADES team (pocket detectors) has discussed
with Wolfram to get access to the ESR tomorrow
(Thu) morning ~8:00 to check their detectors
  
  89   Wed Apr 2 18:22:30 2025 IDGeneralSpill cycle distributionThe spill cycle distributions is absolutely
nuts since beginning of the shift. On the
positive side: we got a lot of spills since
  
ELOG V3.1.5-fc6679b